# Universal framework for anisotropic particles with resonance laws and splitting

### v6

@@ Results
Figure 7: Scattering spectrum and modes/fields near resonance for a biaxial $\alpha$-MoO3 ellipsoid with $(a,b,c)=(1,1.25,1.5)\cdot 11.5\,\mathrm{nm}$ excited by an oscillating dipole located at $\mathbf{r}_{0}=(22,22,22)\,\mathrm{nm}$ with a dipole moment of $\mathbf{P}=(1,1,1)\,\mathrm{A\cdot m}.$. Top: Surface integral over the particle envelope $\int|\mathbf{E}|^{2}da$ calculated in COMSOL and compared to the peaks predicted analytically using $1/|\mathrm{resonance\,condition}|$ with excellent agreement. (a),(b),(c),(d),(e),(f),(g) $|\mathbf{E}|$ for the resonance conditions: $\epsilon_{1x}=\epsilon_{1x,\mathrm{iso}},$ $\epsilon_{1y}=\epsilon_{1y,\mathrm{iso}},$ $\epsilon_{1x}/a^{2}+\epsilon_{1y}/b^{2}=\epsilon_{1xy,\mathrm{iso}}(a^{-2}+b^{-2}),$ $\epsilon_{1x}/a^{2}+\epsilon_{1z}/c^{2}=\epsilon_{1xz,\mathrm{iso}}(a^{-2}+c^{-2}),$ $\epsilon_{1y}/b^{2}+\epsilon_{1z}/c^{2}=\epsilon_{1yz,\mathrm{iso}}(b^{-2}+c^{-2}),$ $\epsilon_{1x}/a^{2}+\epsilon_{1y}/b^{2}=\epsilon_{1xy,\mathrm{iso}}(a^{-2}+b^{-2}),$ $\epsilon_{1x}/a^{2}+\epsilon_{1y}/b^{2}+\epsilon_{1z}/c^{2}=\epsilon_{1xyz,\mathrm{iso}}(a^{-2}+b^{-2}+c^{-2}),$ calculated at $f=28.033,29.467,26.6733,27.467,29.233,29.3,26.6\,\mathrm{(THz)}$ respectively, using COMSOL with very good agreement to the mode fields calculated analytically (same as in Fig. 5, defined in the ellipsoid volume).
@@
Finally, we demonstrate our theory for spheres and ellipsoids composed of hBN and $\alpha-\mathrm{MoO_{3}}$ in full-wave electrodynamic simulations. While our theory applies to any uniaxial and biaxial material, we have chosen these materials as they are currently at the forefront of photonics research and enable to achieve high Q factors. We consider a subwavelength uniaxial hBN sphere with the equal permittivity axes $x,y$ and the permittivity model used in Figs. 2,3 [54]. The sphere has a radius of $11.5\mathrm{nm},$ and we situate an oscillating dipole at $\mathbf{r}_{0}=(22\,\mathrm{nm},0,22\,\mathrm{nm})$ with a dipole moment of $\mathbf{P}=(1,0,1)\,\mathrm{A\cdot m}.$ We perform frequency domain simulations with a frequency sweep from $f=4.3652\cdot 10^{13}\,$(1/s) to $f=4.7626\cdot 10^{13}\,$(1/s) with 304 frequencies using extensive COMSOL simulations. The dipole moment was chosen to be in the direction of $\hat{r}$ since we assume [67, 72, 31], and prove in the SM, that the mode excitation coefficient depends on $\nabla\psi_{l}(\mathbf{r}_{0})\cdot\mathbf{p}$ so that we need to incorporate only $\partial\psi_{l}/\partial r,$ preserving the angular dependence in $\nabla\psi_{l}(\mathbf{r}_{0})\cdot\mathbf{p}$. This, together with the fact that the dipole has only $x,z$ components, ensures that the degenerate modes will add constructively, as is explained in the Methods Section. Note that the coefficient’s dependence on the mode fields at the dipole location, with the rapid decay of the high-order modes away from the particle, reflect the presence of higher spatial frequencies near the dipole, which excite higher-order modes. For a general incoming field, the mode excitation coefficient is given by $\left\langle\psi_{k}|\psi_{0}\right\rangle=\int\theta_{1}\nabla\psi_{k}^{*}\cdot\nabla\psi_{0}d\mathbf{r^{\prime}},$ where $\theta_{1}$ is a window function that equals one in the particle volume, and we utilized a dipole source since a far field excites high-order modes with negligible magnitudes; more details are given in the SM. We then calculate the resonance peak widths analytically (from 1/resonance condition) and from the simulations. To demonstrate the mode-coefficient dependency on the dipole location and orientation, we also perform the same calculation for a dipole oriented along the $x$ axis. In this case, we do not expect the $m=\pm(l-1)$ modes to show in the spectrum since in order to excite them a $z$ component of the dipole location is required. In Fig. 4 (top panel) we present the analytically-predicted peaks in the scattering spectrum obtained from the maxima of $1/|\mathrm{resonance}\,\mathrm{condition}|$ compared to the surface integral of $|\mathbf{E}|^{2}$ over the uniaxial particle envelope calculated in COMSOL with excellent agreement. Importantly, these simulation results for the scattering spectrum confirm our predictions of the resonance splitting (marked by arrows) and mode degeneracy for uniaxial spheres. We also obtained the peak widths analytically from $1/|\mathrm{resonance\,condition}(\omega)|^{2}=\mathrm{max}/2$ and compared them to the ones calculated from the simulations (as can be seen, the width of the peak, which corresponds to $\epsilon_{x}=-1.5,$ is altered due to the peak on the right, and therefore we haven’t analyzed it). Clearly, there is remarkable agreement between them, which showcases our important ability to analytically predict the Q factors given by $\omega/\Delta\omega.$ We then plot for all the observed peaks $|\mathbf{E}|$ inside the particle calculated in COMSOL and obtained analytically from the modes (Methods). While the COMSOL fields include the incoming field from the oscillating dipole, close to a resonance this is typically negligible and indeed we observe very good agreement. Moreover, we plot $\int|\mathbf{E}|^{2}da$ for a dipole located at the same distance as the previous case and oriented along $x$ ($\mathbf{r}_{0}=(1,0,0)$31nm). As we predicted the $m=\pm(l-1)$ modes disappeared from the spectrum, verifying the dependency of the mode excitation coefficient on the dipole location and orientation. This highlights the directional nature of the modes as opposed to isotropic particles. In Fig. 5 we compared the radiation pattern from the full-wave closed-form solution corresponding to the mode $\psi_{2}^{2}+\psi_{2}^{-2}$ with the resonance condition $\epsilon_{1x}=\epsilon_{1y}=-1.5$ outside a uniaxial particle to a full-wave COMSOL simulation of the scattered field of a uniaxial particle with $R=46$ nm and $\epsilon_{1x}=-1.49,\epsilon_{1y}=-1.49,\epsilon_{1z}=1,$ excited by a current source similarly to Fig. 4 (top) in the radiation zone in the radiation zone $(r>2\lambda)$. Note that the COMSOL simulations utilized the axial symmetry and are therefore presented on half a sphere. As can be seen, the radiation-pattern results show very good agreement.
@@
Finally, we simulated a biaxial subwavelength ellipsoid of the most general geometry with $(a,b,c)=(1,1.25,1.5)\cdot 11.5\,\mathrm{nm},$ composed of $\alpha$-MoO3 with the permittivity model in Ref. [54]. The ellipsoid is excited by a dipole located at $\mathbf{r}_{0}=(22\,\mathrm{nm},22\,\mathrm{nm},22\,\mathrm{nm})$ with the dipole moment of $\mathbf{P}=(1,1,1)\,\mathrm{A\cdot m}.$ We performed frequency domain simulations with a frequency sweep between $f=25.5\,\mathrm{(THz)}$ and $f=30.8\,\mathrm{(THz)}$ with 175 frequencies using COMSOL. In Fig. 7 we present the result of the surface integral $\int|\mathbf{E}|^{2}da$ over the particle envelope from the COMSOL simulations and the predicted peaks from the resonance conditions with very good agreement. The predicted peaks were readily calculated by finding the maxima of $1/|\mathrm{resonance}\,\mathrm{condition}|$. We then compared for all the observed resonances $|\mathbf{E}|$ from the simulations to the ones obtained analytically from the modes (equivalent to Fig. 6) with excellent agreement. We note that the isotropic-ellipsoid eigenpermittivities that were calculated to obtain the biaxial ellipsoid resonance permittivity relations are the following: $\epsilon_{\mathrm{iso}\,x}=-1.387,$ $\epsilon_{\mathrm{iso}\,y}=-2.092,$ $\epsilon_{\mathrm{iso}\,xy}=-1.295,$ $\epsilon_{\mathrm{iso}\,yz}=-1.748,$ $\epsilon_{\mathrm{iso}\,xz}=-1.385,$ $\epsilon_{\mathrm{iso}\,xyz}=-1.287.$ As can be seen, for an isotropic ellipsoid with these semi axis ratios $\epsilon_{\mathrm{iso}\,x}\approx\epsilon_{\mathrm{iso}\,xz}$ and $\epsilon_{\mathrm{iso}\,xy}\approx\epsilon_{\mathrm{iso}\,xyz},$ and due to the anisotropic material these resonances are split, as can be seen in Fig. 7. In contrast to spheres, where the degeneracy is enforced by symmetry, here the degeneracy is accidental. Finally, there is a clear physical intuition for the resonance conditions of anisotropic ellipsoids: the resonances of isotropic ellipsoids, on which they are based, tend to favor shorter ellipsoid axes, since they exhibit higher real permittivity values that have smaller imaginary parts of the physical permittivity, leading to stronger and sharper resonances.
Discussion
We derived the modes and eigenpermittivity relations of uniaxial and biaxial particles in the quasistatic regime. We showed that while isotropic spheres have discrete eigenpermittivities, anisotropic spheres exhibit eigenpermittivity sum rules that couple the axial permittivities, leading to resonance splitting. We also identified eigenfrequency degeneracy for uniaxial spheres, which leads to superimposed modes. We generalized our results to uniaxial and biaxial ellipsoids, with eigenpermittivity sum rules that account for both the anisotropic material properties and the ellipsoid’s geometry. Our derived modes and resonance conditions have implications for the thermal emission, enhancement of spontaneous emission, and energy transfer of a single anisotropic particle, as well as for the response of clusters of anisotropic particles [73, 74, 75, 76, 32]. The multispectral behavior of uniaxial and biaxial particles may allow them to function as biomarkers with enhanced specificity of the spectral signature. In addition, the environment-dependent resonance shifts of the particle can be harnessed to measure the biological environment, which may lead to improved cancer detection [59]. Moreover, due to the directional nature of the modes in anisotropic particles, another potential application of our resonators is detection of dark matter [60, 77, 78]. Furthermore, as our derived modes of anisotropic particles correspond to novel radiation patterns, they are expected to have implications for the emission in quantum transitions of atoms in magnetic fields and anisotropic molecules. Future research could focus on analyzing their nonlinear optical response with potential applications in optical computing. Finally, our results apply also to other fields of physics such as heat conduction and quasimagnetostatics [56].
Acknowledgement
I Methods
I.1 Correspondence between the derived modes and the radiation patterns
@@
I.2 Biaxial sphere modes as superpositions of spherical harmonics
@@
I.3 Superimposing the degenerate uniaxial modes for the sphere simulation comparison
I.3.1 Theory

### v7

@@ Methods
Figure 7: Scattering spectrum and modes/fields near resonance for a biaxial $\alpha$-MoO3 ellipsoid with semiaxes $(a,b,c)=(1,1.25,1.5)\cdot 11.5\,\mathrm{nm}$ excited by an oscillating dipole located at $\mathbf{r}_{0}=(22,22,22)\,\mathrm{nm}$ with a dipole moment of $\mathbf{P}=(1,1,1)\,\mathrm{A\cdot m}.$. Top: Surface integral over the particle envelope $\int|\mathbf{E}|^{2}da$ calculated in COMSOL and compared to the peaks predicted analytically using $1/|\mathrm{resonance\,condition}|$ with excellent agreement. (a),(b),(c),(d),(e),(f),(g) $|\mathbf{E}|$ for the resonance conditions: $\epsilon_{1x}=\epsilon_{1x,\mathrm{iso}},$ $\epsilon_{1y}=\epsilon_{1y,\mathrm{iso}},$ $\epsilon_{1x}a^{-2}+\epsilon_{1y}b^{-2}=\epsilon_{1xy,\mathrm{iso}}(a^{-2}+b^{-2}),$ $\epsilon_{1x}a^{-2}+\epsilon_{1z}c^{-2}=\epsilon_{1xz,\mathrm{iso}}(a^{-2}+c^{-2}),$ $\epsilon_{1y}b^{-2}+\epsilon_{1z}c^{-2}=\epsilon_{1yz,\mathrm{iso}}(b^{-2}+c^{-2}),$ $\epsilon_{1x}a^{-2}+\epsilon_{1y}b^{-2}=\epsilon_{1xy,\mathrm{iso}}(a^{-2}+b^{-2}),$ $\epsilon_{1x}a^{-2}+\epsilon_{1y}b^{-2}+\epsilon_{1z}c^{-2}=\epsilon_{1xyz,\mathrm{iso}}(a^{-2}+b^{-2}+c^{-2}),$ calculated at $f=28.033,29.467,26.6733,27.467,29.233,29.3,26.6\,\mathrm{(THz)}$ respectively, using COMSOL with very good agreement to the mode fields calculated analytically (same as in Fig. 6, defined in the ellipsoid volume).
@@
Finally, we demonstrate our theory for spheres and ellipsoids composed of hBN and $\alpha-\mathrm{MoO_{3}}$ in full-wave electrodynamic simulations. While our theory applies to any uniaxial and biaxial material, we have chosen these materials as they are currently at the forefront of photonics research and enable to achieve high Q factors. We consider a subwavelength uniaxial hBN sphere with the equal permittivity axes $x,y$ and the permittivity model used in Figs. 2,3 [54]. The sphere has a radius of $11.5\mathrm{nm},$ and we situate an oscillating dipole at $\mathbf{r}_{0}=(22\,\mathrm{nm},0,22\,\mathrm{nm})$ with a dipole moment of $\mathbf{P}=(1,0,1)\,\mathrm{A\cdot m}.$ We perform frequency domain simulations with a frequency sweep from $f=4.3652\cdot 10^{13}\,$(1/s) to $f=4.7626\cdot 10^{13}\,$(1/s) with 304 frequencies using extensive COMSOL simulations. The dipole moment was chosen to be in the direction of $\hat{r}$ since we assume [67, 78, 32], and prove in the SM, that the mode excitation coefficient depends on $\nabla\psi_{l}(\mathbf{r}_{0})\cdot\mathbf{p}$ so that we need to incorporate only $\partial\psi_{l}/\partial r,$ preserving the angular dependence in $\nabla\psi_{l}(\mathbf{r}_{0})\cdot\mathbf{p}$. This, together with the fact that the dipole has only $x,z$ components, ensures that the degenerate modes will add constructively, as is explained in Methods D. Note that the coefficient’s dependence on the mode fields at the dipole location, with the rapid decay of the high-order modes away from the particle, reflect the presence of higher spatial frequencies near the dipole, which excite higher-order modes. For a general incoming field, the mode excitation coefficient is given by $\left\langle\psi_{k}|\psi_{0}\right\rangle=\int\theta_{1}\nabla\psi_{k}^{*}\cdot\nabla\psi_{0}d\mathbf{r^{\prime}},$ where $\theta_{1}$ is a window function that equals one in the particle volume, and we utilized a dipole source since a far field excites high-order modes with negligible magnitudes; more details are given in the SM. We then calculate the resonance peak widths analytically (from 1/resonance condition) and from the simulations. To demonstrate the mode-coefficient dependency on the dipole location and orientation, we also perform the same calculation for a dipole oriented along the $x$ axis. In this case, we do not expect the $m=\pm(l-1)$ modes to show in the spectrum since in order to excite them a $z$ component of the dipole location is required. In Fig. 4 (top panel) we present the analytically-predicted peaks in the scattering spectrum obtained from the maxima of $1/|\mathrm{resonance}\,\mathrm{condition}|$ compared to the surface integral of $|\mathbf{E}|^{2}$ over the uniaxial particle envelope calculated in COMSOL with excellent agreement. Importantly, these simulation results for the scattering spectrum confirm our predictions of the resonance splitting (marked by arrows) and mode degeneracy for uniaxial spheres. We also obtained the peak widths analytically from $1/|\mathrm{resonance\,condition}(\omega)|^{2}=\mathrm{max}/2$ and compared them to the ones calculated from the simulations (as can be seen, the width of the peak, which corresponds to $\epsilon_{x}=-1.5,$ is altered due to the peak on the right, and therefore we haven’t analyzed it). Clearly, there is remarkable agreement between them, which showcases our important ability to analytically predict the Q factors given by $\omega/\Delta\omega.$ We then plot for all the observed peaks $|\mathbf{E}|$ inside the particle calculated in COMSOL and obtained analytically from the modes (Methods). While the COMSOL fields include the incoming field from the oscillating dipole, close to a resonance this is typically negligible and indeed we observe very good agreement. Moreover, we plot $\int|\mathbf{E}|^{2}da$ for a dipole located at the same distance as the previous case and oriented along $x$ ($\mathbf{r}_{0}=(1,0,0)$31nm). As we predicted the $m=\pm(l-1)$ modes disappeared from the spectrum, verifying the dependency of the mode excitation coefficient on the dipole location and orientation. This highlights the directional nature of the modes as opposed to isotropic particles. In Fig. 5 we compared the radiation pattern from the full-wave closed-form solution corresponding to the mode $\psi_{2}^{2}+\psi_{2}^{-2}$ with the resonance condition $\epsilon_{1x}=\epsilon_{1y}=-1.5$ outside a uniaxial particle to a full-wave COMSOL simulation of the scattered field of a uniaxial particle with $R=46$ nm and $\epsilon_{1x}=-1.49,\epsilon_{1y}=-1.49,\epsilon_{1z}=1,$ excited by a current source similarly to Fig. 4 (top) in the radiation zone in the radiation zone $(r>2\lambda)$. Note that the COMSOL simulations utilized the axial symmetry and are therefore presented on half a sphere. As can be seen, the radiation-pattern results show very good agreement.
@@
Finally, we simulated a biaxial subwavelength ellipsoid of the most general geometry with semiaxes $(a,b,c)=(1,1.25,1.5)\cdot 11.5\,\mathrm{nm},$ composed of $\alpha$-MoO3 with the permittivity model in Ref. [54]. The ellipsoid is excited by a dipole located at $\mathbf{r}_{0}=(22\,\mathrm{nm},22\,\mathrm{nm},22\,\mathrm{nm})$ with the dipole moment of $\mathbf{P}=(1,1,1)\,\mathrm{A\cdot m}.$ We performed frequency domain simulations with a frequency sweep between $f=25.5\,\mathrm{(THz)}$ and $f=30.8\,\mathrm{(THz)}$ with 175 frequencies using COMSOL. In Fig. 7 we present the result of the surface integral $\int|\mathbf{E}|^{2}da$ over the particle envelope from the COMSOL simulations and the predicted peaks from the resonance conditions with very good agreement. The predicted peaks were readily calculated by finding the maxima of $1/|\mathrm{resonance}\,\mathrm{condition}|$. We then compared for all the observed resonances $|\mathbf{E}|$ from the simulations to the ones obtained analytically from the modes (equivalent to Fig. 6) with excellent agreement. We note that the isotropic-ellipsoid eigenpermittivities that were calculated to obtain the biaxial ellipsoid resonance permittivity relations are the following: $\epsilon_{\mathrm{iso}\,x}=-1.387,$ $\epsilon_{\mathrm{iso}\,y}=-2.092,$ $\epsilon_{\mathrm{iso}\,xy}=-1.295,$ $\epsilon_{\mathrm{iso}\,yz}=-1.748,$ $\epsilon_{\mathrm{iso}\,xz}=-1.385,$ $\epsilon_{\mathrm{iso}\,xyz}=-1.287.$ As can be seen, for an isotropic ellipsoid with these semi axis ratios $\epsilon_{\mathrm{iso}\,x}\approx\epsilon_{\mathrm{iso}\,xz}$ and $\epsilon_{\mathrm{iso}\,xy}\approx\epsilon_{\mathrm{iso}\,xyz},$ and due to the anisotropic material these resonances are split, as can be seen in Fig. 7. In contrast to spheres, where the degeneracy is enforced by symmetry, here the degeneracy is accidental. Finally, there is a clear physical intuition for the resonance conditions of anisotropic ellipsoids: the resonances of isotropic ellipsoids, on which they are based, tend to favor shorter ellipsoid axes, since they exhibit higher real permittivity values that have smaller imaginary parts of the physical permittivity, leading to stronger and sharper resonances.
.2 Correspondence between the derived modes and the radiation patterns
@@
.3 Biaxial sphere modes as superpositions of spherical harmonics
@@
.4 Superimposing the degenerate uniaxial modes for the sphere simulation comparison
.4.1 Theory